# Lessons learned from field demonstrations of model predictive control and reinforcement learning for residential and commercial HVAC: A review

### v1

@@ Acknowledgments
We sincerely thank the following researchers for responding to our requests for further information about their field demonstrations: José Domingo Álvarez Hervás (University of Almería), David Blum (Lawrence Berkeley National Laboratory [LBNL]), Hanmin Cai (Swiss Federal Laboratories for Materials Science and Technology), Roel De Coninck (dnergy), Ján Drgoňa (Johns Hopkins University, Pacific Northwest National Laboratory), Svenne Freund (ROM Technik), Jessica Granderson (LBNL), Sang woo Ham (LBNL), Lieve Helsen (KU Leuven), Hao Huang (Frasers Property Limited), Yilin Jiang (University of Oklahoma, LBNL), Donghun Kim (LBNL), Michael Dahl Knudsen (Aarhus University), David Lindelöf (Google), Iakovos Michailidis (The Centre for Research & Technology Hellas), Avisek Naug (Hewlett Packard Labs), Zheng O’Neill (Texas A&M University), Marco Pritoni (LBNL), Etienne Saloux (CanmetENERGY), Jicheng Shi (École Polytechnique Fédérale de Lausanne), Jan Široký (Energy Twin), Phillip Stoffel (E.ON Energy Research Center - RWTH Aachen University), Jacopo Vivian (University of Padua), Xuezheng Wang (Syracuse University), Sicheng Zhan (Massachusetts Institute of Technology), and Kun Zhang (École de Technologie Supérieure Montréal, LBNL). We gratefully acknowledge scholarship support from the Onassis Foundation (E.N. Pergantis; P. Papageorgiou) and Graduate Research Fellowships Program support from the National Science Foundation (L.D. Reyes Premer). Thanks as well to Ettore Zanetti (LBNL) for helpful discussion.
Appendix A Field demonstration summary tables
@@
# Study (Year) Control Method Location Building Description HVAC System Test Duration (days) Objective(s) Reported Outcome(s)
@@
MPC Denmark four single-family detached houses radiant system with ground-to-water heat pumps, floor heating, and thermal energy storage (TES) in concrete floors 152 minimize cost 9.17% cost savings vs. measured baseline
@@
sustainable house, visitor location hybrid radiant floor heating and forced air cooling with fixed settings 10 minimize energy 17.8–30.1% energy savings vs. simulated baseline
@@
MPC Switzerland and Germany eight single-family houses and two apartments radiant system with water distribution; sources include wood pellets, heat pumps, gas, and oil 27–138 minimize energy 3.87–60.4% energy savings vs. measured baseline
@@
RL Leuven, Belgium living lab setup featuring a test room forced air system integrated with photovoltaic (PV) integration 3 maximize solar self-utilization reduced PV peak power injection and synchronized cooling with PV generation vs. measured baseline
@@
MPC Roskilde, Denmark single family detached house radiant water-based heating with electro-valve radiators 10 minimize energy, minimize cost and shift loads to off-peak periods -3.2–5% energy savings, 3–11.3% cost savings, and effective load shifting during off-peak hours vs. simulated baseline
@@
MPC Vaughan, ON, Canada four floors (including basement) in the TRCA-ASH, with an adjacent in-law suite hybrid system with radiant heating, forced-air cooling, and a ground source heat pump 22 minimize cost and energy 16–142.3% cost and -1.5–114.6% energy vs. simulated baseline
@@
MPC Dübendorf, Switzerland Urban Mining and Recycling (UMAR) unit, a residential module in Empa’s NEST demonstrator air-to-water heat pump with hydronic radiant ceilings and room-specific valves 6 minimize energy 24.9% vs. measured baseline
@@
MPC near Amstelveen, Netherlands detached low-energy house PVT to HP system supplying SH and DHW tanks with hydronic floor heating 1 minimize cost 12–25% vs. measured baseline
@@
RL Knoxville, TN, USA detached residential house complying with IECC 2006 standards two-zone air-to-air AC with two-stage compressor and variable-speed fan 5 minimize cost 11–21% vs. simulated baseline
@@
MPC Trondheim, Norway research Centre on Zero Emission Buildings (ZEB) constant tank heating, ventilation, and hydronic floor heating 14 minimize cost 22.5% vs. simulated baseline
@@
RL Dübendorf, Switzerland UMAR unit, a residential module in Empa’s NEST demonstrator radiant heating with a heat pump and emulated EV integration 5 minimize energy 27% vs. measured baseline
@@
MPC Dübendorf, Switzerland UMAR unit, a residential module in Empa’s NEST demonstrator hydronic loop with heat pumps, variable-speed pump, room valves, and ceiling radiant systems 35–41 minimize energy 26–49% vs. measured baseline
@@
MPC Piacenza, Italy lightweight building lab air-to-water heat pump, fan coils, air/air system, water heater, air extractors, simulated PV panels 5 minimize cost 10–17% vs. measured baseline
@@
MPC Ottawa, ON, Canada CHEeR house, a two-story research facility at Carleton University hydronic floor system for heating and cooling 14 minimize energy 40.6% vs. measured baseline
@@
MPC Shenzhen, China two-room apartment with living room mini-split air conditioners, fixed-speed, single-stage 12 minimize cost 22.1–26.8% vs. measured baseline
@@
MPC Shawinigan, QC, Canada two-story detached house with a basement radiant heating with electric baseboards and smart thermostats. No active cooling. 3 minimize energy and cost 71% energy and 55% cost vs. measured baseline
@@
MPC Zealand, Denmark detached residential building classified as a low-energy house (BR2020) variable-speed air-to-water heat pump for hydronic space heating and domestic hot water 10–50 minimize cost 2.3–32.8% vs. measured baseline
@@
MPC Ottawa, ON, Canada detached residential building classified as a low-energy house (CHEeR house, a two-story research facility at Carleton University) hydronic floor system for heating and cooling 182 minimize energy runtime optimized; savings inferred vs. measured baseline
19 [124]
2023
MPC Dübendorf, Switzerland UMAR unit in the NEST building at Empa, a three-room residential module hydronic radiant ceiling panels, heating controlled via valve regulation and pulse-width modulation 5 minimize energy savings not reported; validated experimentally vs. simulated baseline
@@
MPC Norman, OK and Miami, FL, USA 9 single-family homes central forced-air HVAC systems with AC units designed for residential cooling 2–24 minimize energy and cost -29.76–91.05% energy and -13.49–87.06% cost vs. measured baseline
@@
MPC West Lafayette, IN, USA DC Nanogrid house, a detached single-family home forced-air, variable-speed air-to-air heat pump with staged electric resistance backup heating 33 minimize energy, cost, and peak demands 19% avg energy and 28% avg cost vs. measured baseline
@@
MPC Reduit, Mauritius, Africa single-story, two-zone concrete building designed for controlled experiments in a tropical climate each room with a portable AC, fan, air extractors, forced-air system, adjustable speeds $<1$ minimize temperature tracking error improved temperature tracking and efficiency vs. measured baseline
@@
MPC West Lafayette, IN, USA DC Nanogrid house, a detached single-family home forced-air, variable-speed air-to-air heat pump with staged electric resistance backup heating 38 minimize cost and peak demands 19% avg energy and 28% avg cost vs. measured baseline
@@
MPC West Lafayette, IN, USA DC Nanogrid house, a detached single-family home forced air, variable-speed air-to-air heat pump with staged electric resistance backup heating 31 enable electrification within 100A panels cost-effectively maintained peak current within 100 A, avoided costly upgrades, and enabled additional EV charging vs. measured baseline
@@
# Study (Year) Control Method Location Building Description HVAC System Test Duration (days) Objective(s) Reported Outcome(s)
@@
MPC Ankeny, IA, US 8 office rooms at Energy Resource Station (ERS) forced-air HVAC with fixed-speed operation, air-cooled chillers, ice storage, AHUs, and auxiliaries 4 minimize cost experimental savings negligible; simulated savings 17–27% vs. simulated baseline
@@
RL Ankeny, IA, US 8 office rooms in ERS forced-air HVAC with fixed-speed operation, air-cooled chillers, ice storage, AHUs, and auxiliaries 6 minimize cost 8.3% vs. simulated baseline
@@
MPC Chania, Crete, Greece laboratory at the Technical University of Crete forced air system. heating and cooling via air conditioning. 3 minimize energy experimental savings unreported; vs. measured baseline
@@
MPC Garston, England Building Research Establishment (BRE), with spaces including offices, meeting rooms, and lecture theaters low-pressure radiant heating with thermostatic valves 45 minimize energy experimental savings not reported; simulated savings 9–32.1% vs. simulated baseline
@@
MPC Prague, Czech Republic Czech Technical University building: seven control blocks hydronic ceiling radiant heating: embedded beams, heat exchanger, valve control 45 minimize energy and cost 20.5–29% energy and 29% cost vs. measured baseline
@@
MPC Prague, Czech Republic Czech Technical University building: seven control blocks hydronic ceiling radiant heating: embedded beams, heat exchanger, valve control 7–49 minimize energy 16.24–28.74% energy vs. measured baseline
@@
MPC Almería, Spain an office room within a bioclimatic research building at the University of Almería forced-air HVAC system with adjustable fan speed and water flow not reported minimize energy not reported vs. simulated baseline
@@
MPC Berkeley, CA, US a ground-floor computer laboratory in a UC Berkeley engineering building an electric single-stage forced-air heat pump air conditioner with a single-speed operation 3 minimize energy 28–66% energy vs. simulated baseline
@@
MPC Merced, CA, US campus building in University of California, Merced centralized HVAC system with chillers, thermal energy storage, and distributed water loops 5 minimize cost and maximize efficiency (COP) \$1,280 daily cost savings and 19.1% COP improvement vs. simulated baseline
@@
MPC Atlanta, GA and Phoenix, AZ, US test chambers: a typical office zone with an exterior wall and a climate chamber simulating outdoor conditions chilled water hydronic loop with variable-speed control 7 minimize energy 19–25% energy vs. measured baseline
@@
MPC Algarve, Portugal four rooms at the University of Algarve: a computer classroom, adjacent spaces, and a larger room with two units forced-air VRF systems with outdoor and internal units in each room 1–2 minimize energy 57% energy vs. measured baseline
@@
MPC Champaign, IL, US office building at US Army Corps of Engineers Construction Engineering Research Laboratory (CERL) forced-air variable-volume system with a centralized dual-duct multi-zone AHU 5 minimize energy 60–85% energy vs. measured baseline
@@
MPC Newcastle and Melbourne, Australia three-story office buildings forced-air system with AHUs supplied by chillers, a CHP system, and boilers 5–25 minimize energy and cost 29–32% energy and 16.9% cost vs. measured baseline
@@
MPC Almería, Spain an office room within a bioclimatic research building at the University of Almería forced-air HVAC system with adjustable fan speed and water flow 15 minimize energy 53% energy vs. measured baseline
@@
MPC Hasselt, Belgium Hollandsch Huys office building; three floors, excluding underground garage and under-roof apartment hydronic ground-coupled heat pumps; thermally activated building systems (TABS), AHUs, ground-floor heating; seasonal thermal storage 25 minimize energy 11.37–22.77% energy vs. measured baseline
@@
MPC Berkeley, CA, US Sutardja Dai Hall, a modern building on the UC Berkeley campus, with classrooms, labs, offices forced-air variable-speed fans and VFDs $<1$ minimize energy and maximize flexibility 25% variation in fan power consumption and up to 20% downward flexibility vs. measured baseline
@@
MPC Graz, Austria research facility, designed to simulate real-world conditions in a controlled environment forced-air system, including fans, heating and cooling coils, steam humidifiers, air dampers, and filters $<1$ minimize energy 44% energy vs. measured baseline
@@
MPC Stockholm, Sweden ground-floor laboratory in the Q-building on the KTH Royal Institute of Technology campus ventilation with central fan; heating from district radiators; air inlet with cooling coil 2 minimize energy not reported vs. measured baseline
@@
MPC Stockholm, Sweden ground-floor laboratory in the Q-building on the KTH Royal Institute of Technology campus radiator system with district-heated water for indoor heating 1 minimize energy 31–33% energy vs. measured baseline
@@
MPC Gainesville, FL, US an office room in Pugh Hall, University of Florida forced-air system with one AHU and VAV for zone control
1 minimize energy 40% energy vs. measured baseline
@@
MPC Ankeny, IA, US 4 office rooms in ERS one AHU connected to multiple VAVs, utilizing a forced-air system
1 minimize energy 2.23–17.24% energy vs. simulated baseline
@@
MPC Adelaide, Australia check-in hall of T1 building at Adelaide Airport, South Australia; four thermal zones at the perimeter areas of level 2 forced-air cooling with chillers, variable-speed pumps, and AHUs 1 minimize cost 13% cost vs. measured baseline
@@
MPC Philadelphia, PA, US one-third of the north side, four floors, in Building 101, a medium-sized office at the Philadelphia Navy Yard forced-air system with three AHUs, VAV boxes, DX cooling, and gas boiler heating; focus on AHU3 20 minimize energy 33% energy vs. measured baseline
@@
MPC Knoxville, TN, US basketball court within a gym affiliated with ORNL forced-air system with four two-stage RTUs
4 minimize energy and peak demand 8% energy and 43% peak demand vs. measured baseline
@@
MPC Allschwil, Switzerland office building, with five office floors and a ground floor housing a kitchen and restaurant hybrid system with forced-air AHU, radiant TABS, gas boiler heating, dry-mode cooling tower, and corner office radiators 203 minimize cost experimental savings unreported; simulated savings 17% vs. simulated baseline
@@
MPC Brussels, Belgium medium-sized office building with two floors hybrid HVAC system with air-source heat pumps, boiler, and hydronic distribution 15 minimize energy and cost 20–30% energy and 34–40% cost vs. measured baseline
@@
MPC Faro, Portugal three rooms in Building 7, Gambelas campus, University of Algarve forced-air VRF system with an air-cooled compressor and ceiling-concealed units $<1$ minimize energy 50–51% energy vs. measured baseline
@@
RL Genk, Belgium research space designed for real-world energy testing two air conditioners with basic monitoring and control setup 16 minimize cost exp. savings unreported (no benchmark); sim. performance within 90% of mathematical optimum vs. simulated baseline
@@
MPC Lausanne, Switzerland four office rooms in the Automation Control Lab at École Polytechnique Fédérale de Lausanne (EPFL) radiant heating system with commercial electric heater, original hot water-based radiators switched off $<1$ frequency regulation evaluation improved frequency regulation vs. simulated baseline
@@
MPC Berkeley, CA, US thermally identical test cells in FLEXLAB at Lawrence Berkeley National Laboratory (LBNL) forced-air HVAC systems with variable speed and dedicated AHUs 1 frequency regulation evaluation improved frequency regulation vs. measured baseline
@@
MPC Lausanne, Switzerland four-decade-old office building with five distinct office spaces, including shared and individual rooms electric heaters with direct modulation and room-specific control 1 frequency regulation evaluation improved frequency regulation vs. simulated baseline
@@
Granada, Spain three-story school hydronic heating system with a biomass boiler serving separate circuits for main building and library 43 minimize energy 33% energy vs. measured baseline
@@
MPC Neuchâtel and Winterthur, Switzerland a conference room in the Swiss Center for Electronics and Microtechnology (CSEM) and an office building forced-air HVAC systems with AHUs, standard radiators, and a rotating heat exchanger 10–38 minimize cost 21–23% cost vs. measured baseline
@@
MPC Halifax, NS, Canada LEED Gold academic building at Dalhousie University, spanning five floors forced-air HVAC system with VAV reheat, heat pumps, district heating, and rooftop cooling tower 113 minimize energy 29–63% energy vs. measured baseline
@@
MPC Salzburg, Austria two floors with 500 rooms in a five-floor office building at the University of Salzburg hybrid HVAC system with radiant TABS and forced air Fan Coils for heating and cooling 49 minimize energy experimental savings unreported; simulated savings 31–36% vs. simulated baseline
@@
MPC Sichuan Province, China shopping mall with six above-ground floors and two below-ground levels forced-air system with centrifugal chillers, variable speed drivers, AHUs, and fan coils 1 minimize energy 16% energy vs. measured baseline
@@
RL Aachen, Germany three conference rooms at the E.ON Energy Research Center, RWTH Aachen University hybrid forced air and radiant system with concrete core activation and air chillers 5 minimize energy 34.7% energy vs. measured baseline

### v6

@@ Acknowledgments
We sincerely thank the following researchers for responding to our requests for further information about their field demonstrations: José Domingo Álvarez Hervás (University of Almería), David Blum (Lawrence Berkeley National Laboratory [LBNL]), Hanmin Cai (Swiss Federal Laboratories for Materials Science and Technology), Roel De Coninck (dnergy), Ján Drgoňa (Johns Hopkins University), Svenne Freund (ROM Technik), Jessica Granderson (LBNL), Sang woo Ham (LBNL), Lieve Helsen (KU Leuven), Hao Huang (Frasers Property Limited), Yilin Jiang (Pacific Northwest National Laboratory), Donghun Kim (LBNL), Michael Dahl Knudsen (Aarhus University), David Lindelöf (Google), Iakovos Michailidis (The Centre for Research & Technology Hellas), Avisek Naug (Hewlett Packard Labs), Zheng O’Neill (Texas A&M University), Marco Pritoni (LBNL), Etienne Saloux (CanmetENERGY), Jicheng Shi (École Polytechnique Fédérale de Lausanne), Jan Široký (Energy Twin), Phillip Stoffel (E.ON Energy Research Center - RWTH Aachen University), Jacopo Vivian (University of Padua), Xuezheng Wang (Syracuse University), Sicheng Zhan (Massachusetts Institute of Technology), and Kun Zhang (École de Technologie Supérieure Montréal). Thanks as well to Ettore Zanetti (LBNL) for helpful discussion. We gratefully acknowledge scholarship support from The American Society of Heating, Refrigerating and Air-Conditioning Engineers Grant-In-Aid Award Program (A.J. Khabbazi; E.N. Pergantis; L.D. Reyes Premer), scholarship support from The Alexander S. Onassis Foundation (E.N. Pergantis; P. Papageorgiou), and Graduate Research Fellowships Program support from the U.S. National Science Foundation (L.D. Reyes Premer).
Appendix A Nomenclature
AHU air handling unit
DeePC data-enabled predictive control
DER distributed energy resource
HP heat pump
HVAC heating, ventilation, and air-conditioning
MPC model predictive control
OCC occupant-centric control
PV photovoltaics
RL reinforcement learning
TABS thermally-activate building system
VAV variable air volume
VRF variable refrigerant flow
Appendix B Field demonstration summary tables
@@
# Study (Year) Control Method Location Building Description HVAC System Test Days Objective(s) Reported Outcome(s)
vs. Benchmarking Method
@@
MPC Denmark four single-family detached houses radiant floor heating with ground-to-water heat pumps (HPs) 152 minimize cost 9.2% cost savings vs. measured baseline
@@
a sustainable house, visitor location radiant floor heating and forced-air cooling with fixed settings 10 minimize energy 17.8–30.1% energy savings vs. simulated baseline
@@
MPC Switzerland and Germany eight single-family houses and two apartments radiant with water distribution; sources include wood pellets, HPs, gas, and oil 27–138 minimize energy 3.9–60.4% energy savings vs. measured baseline
@@
RL Leuven, Belgium living lab setup featuring a test room forced-air integrated with PV system 3 maximize solar self-utilization reduced PV peak power injection and synchronized cooling with PV generation vs. measured baseline
@@
MPC Roskilde, Denmark single family detached house radiant water-based heating with electro-valve radiators 10 minimize energy and cost; shift loads to off-peak periods -3.2–5% energy savings, 3–11.3% cost savings, and effective load shift during off-peak hours vs. simulated baseline
@@
MPC Vaughan, ON, Canada four-story sustainable house, including a basement, with an adjacent in-law suite hybrid system with radiant heating, forced-air cooling, and a ground source HP with a buffer tank 22 minimize energy and cost -4.1–52.2% energy savings and 8.5–56.1% cost savings vs. simulated baseline
@@
MPC Dübendorf, Switzerland residential module in a sustainable demonstrator building air-to-water HP with hydronic radiant ceilings and room-specific valves 6 minimize energy 24.9% energy savings vs. measured baseline
@@
MPC near Amstelveen, Netherlands detached four-story low-energy house hydronic floor heating supplied by an HP powered by PV thermal panels 1 minimize cost 12–15% cost savings vs. measured baseline
@@
RL Knoxville, TN, USA detached energy-efficient house two-zone air-to-air conditioning with two-stage compressor and variable-speed fan 5 minimize cost 11–21% cost savings vs. simulated baseline
@@
MPC Trondheim, Norway test house at a university hydronic floor heating from a resistive-heated tank, with mechanical ventilation 14 minimize cost 22.5% cost savings vs. simulated baseline
@@
RL Dübendorf, Switzerland residential module in a sustainable demonstrator building radiant heating with HP and emulated electric vehicle integration 5 minimize energy 27% energy savings vs. measured baseline
@@
MPC Dübendorf, Switzerland residential module in a sustainable demonstrator building hydronic loop with HPs, variable-speed pump, room valves, and ceiling radiant systems 35–41 minimize energy 26–49% energy savings vs. measured baseline
@@
MPC Piacenza, Italy laboratory building air-to-water HP with fan coils, air extractors, and simulated PV panels 5 minimize cost 10–17% cost savings vs. measured baseline
@@
MPC Ottawa, ON, Canada two-story research facility within a university hydronic floor system for heating and cooling 13 minimize energy 40.6% energy savings vs. measured baseline
@@
MPC Shenzhen, China two-room apartment with living room forced-air with fixed-speed, single-stage mini-split air conditioners 12 minimize cost 22.1–26.8% cost savings vs. measured baseline
@@
MPC Shawinigan, QC, Canada two-story detached house with a basement radiant heating with electric baseboards and smart thermostats 3 minimize cost and shift load 55% cost savings and 71% high-price energy reduction vs. measured baseline
@@
MPC Zealand, Denmark detached low-energy house hydronic variable-speed air-to-water HP for space heating and domestic hot water 10–50 minimize cost 2.3–32.8% cost savings vs. measured baseline
@@
MPC Ottawa, ON, Canada two-story research facility within a university hydronic floor system for heating and cooling 182 minimize energy runtime optimized; savings inferred vs. measured baseline
@@
MPC
Dübendorf, Switzerland residential module in a sustainable demonstrator building hydronic loop with HPs, variable-speed pump, room valves, and ceiling radiant systems 5 minimize energy experimental savings unreported; simulated energy savings 12–27% vs. simulated baseline
@@
MPC Norman, OK, and Miami, FL, USA nine single-family houses central forced-air systems with ACs designed for cooling 2–24 minimize energy and cost -29.8–91.1% energy savings and -13.5–87.1% cost savings vs. measured baseline
@@
MPC West Lafayette, IN, USA detached full-electric single-family home forced-air variable-speed air-to-air HP with staged electric resistance backup heating 33 minimize energy and cost 19% energy savings and 28% cost savings vs. measured baseline
@@
MPC Reduit, Mauritius, Africa single-story, two-zone concrete building for tropical climate testing forced-air with a portable AC, electric fan, and air extractors, with adjustable speed settings $<1$ minimize temperature tracking error improved temperature tracking and efficiency vs. measured baseline
23 [129]
2024
MPC West Lafayette, IN, USA detached full-electric single-family home forced-air variable-speed air-to-air HP with staged electric resistance backup heating 38 minimize cost -5–32.5% cost savings vs. measured baseline
@@
MPC West Lafayette, IN, USA detached full-electric single-family home forced-air variable-speed air-to-air HP with staged electric resistance backup heating 31 whole home hard current draw constraint maintained peak current draw below 100 A (no benchmark)
@@
# Study (Year) Control Method Location Building Description HVAC System Test Days Objective(s) Reported Outcome(s)
vs. Benchmarking Method
@@
MPC Ankeny, IA, US eight office rooms at a test facility forced-air fixed-speed with air-cooled chillers, ice storage, and air handling units (AHUs) 4 minimize cost experimental savings negligible; simulated cost savings 17–27% vs. simulated baseline
@@
RL Ankeny, IA, US eight office rooms at a test facility forced-air fixed-speed with air-cooled chillers, ice storage, and AHUs 6 minimize cost 8.3% costing savings vs. simulated baseline
@@
MPC Chania, Crete, Greece laboratory in a university building forced-air heating and cooling via split-type air conditioning 3 minimize energy experimental savings unreported vs. measured baseline
@@
MPC Garston, England office building with office, meeting rooms, and lecture theaters radiant heating with thermostatic valves 45 minimize energy experimental savings unreported; simulated energy savings 9–32.1% vs. simulated baseline
@@
MPC Prague, Czech Republic university building with seven control blocks hydronic ceiling radiant heating with embedded beams, heat exchanger, and valve control 45 minimize energy and cost 20.5–29% energy savings and 29% cost savings vs. measured baseline
@@
MPC Prague, Czech Republic university building with seven control blocks hydronic ceiling radiant heating with embedded beams, heat exchanger, and valve control 7–49 minimize energy 16.24–28.74% energy savings vs. measured baseline
@@
MPC Almería, Spain office space in a bioclimatic research facility forced-air with variable fan speed and chilled water coils NR [^4] Not reported minimize energy experimental savings unreported vs. simulated baseline
@@
MPC Berkeley, CA, US ground-floor university building forced-air, single-stage HP with single-speed operation 3 minimize energy 28–66% energy savings vs. simulated baseline
@@
MPC Merced, CA, US university building forced-air centralized system with chillers, thermal energy storage, and distribution loops 5 minimize cost and maximize coefficient of performance \$1,280 daily cost savings and 19.1% coefficient of performance improvement vs. measured baseline
@@
MPC Atlanta, GA and Phoenix, AZ, US test chamber simulating office conditions with another climate chamber thermally-activate building system (TABS) with radiant cooling by low-lift chilled water 7 minimize energy 19–25% energy savings vs. measured baseline
@@
MPC Algarve, Portugal four rooms in a university building forced-air variable refrigerant flow (VRF) systems with ducted indoor and outdoor units per zone 1–2 minimize energy 41–77% energy savings vs. measured baseline
@@
MPC Champaign, IL, US office building forced-air with variable air volume (VAV) dual-duct multi-zone AHU 5 minimize energy 60–85% energy savings vs. measured baseline
@@
MPC Newcastle and Melbourne, Australia three-story office buildings forced-air with AHUs, chillers, combined heat and power, and boilers 5–25 minimize energy and cost 14–32% energy savings and 16.9% cost savings vs. measured baseline
@@
MPC Almería, Spain office space in a bioclimatic research facility forced-air with variable fan speed and chilled water coils 15 minimize energy 53% energy savings vs. measured baseline
@@
MPC Hasselt, Belgium three-floor office building, underground garage and under-roof apartment excluded hydronic ground-coupled HPs with TABS, AHUs, floor heating, and seasonal thermal storage 25 minimize energy 11.4–22.8% energy savings vs. measured baseline
@@
MPC Berkeley, CA, US university building forced-air variable-speed fans and variable frequency drives $<1$ maximize flexibility 25% fan power consumption variation and up to 20% downward flexibility vs. measured baseline
@@
MPC Graz, Austria research facility simulating real-world conditions forced-air with fans, coils, humidifiers, dampers, and filters $<1$ minimize energy 44% energy savings vs. measured baseline
@@
MPC Stockholm, Sweden ground-floor university laboratory forced-air ventilation, district radiator heating, and inlet coil cooling 2 minimize energy experimental savings unreported vs. measured baseline
@@
MPC Stockholm, Sweden ground-floor university laboratory radiant heating via district-heated water radiators 1 minimize energy 31–33% energy savings vs. measured baseline
@@
MPC Gainesville, FL, US office room in a university building forced-air via AHU, with VAV terminal and reheat for zone control 1 minimize energy 40% energy savings vs. measured baseline
@@
MPC Ankeny, IA, US four office rooms at a test facility forced-air with one AHU connected to multiple VAVs 1 minimize energy 2.2–17.2% energy savings vs. simulated baseline
@@
MPC Adelaide, Australia one zone in the check-in hall of an airport forced-air cooling with chillers, AHUs, and variable-speed pumps 1 minimize cost 13% cost savings vs. measured baseline
@@
MPC Philadelphia, PA, US one-third of a medium-sized four-story office building forced-air with three AHUs, VAVs with direct expansion cooling and gas boiler heating 20 minimize energy 33% energy savings vs. measured baseline
@@
MPC Knoxville, TN, US basketball court within a gym forced-air with four two-stage rooftop units (RTUs) 4 minimize energy and peak demand 8% energy savings and 43% peak demand reduction vs. measured baseline
@@
MPC Allschwil, Switzerland office building with five office floors and ground-floor kitchen and restaurant hybrid system with AHU, TABS, gas boiler, dry cooling tower, and corner radiators 10–98 minimize cost experimental savings unreported; simulated cost savings 17% vs. simulated baseline
@@
MPC Brussels, Belgium two-story office building hybrid system with air-source HPs, boiler, and hydronic loop 15 minimize energy and cost 20–30% energy savings and 34–40% cost savings vs. measured baseline
@@
MPC Faro, Portugal three rooms in a university building forced-air VRF system with air-cooled outdoor unit and ceiling-concealed indoor units $<1$ minimize energy 50–51% energy savings vs. measured baseline
@@
RL Genk, Belgium laboratory space for energy testing ductless air conditioner units with basic monitoring 16 minimize cost experimental savings unreported (no benchmark); simulation performance within 90% of mathematical optimum vs. simulated baseline
@@
MPC Lausanne, Switzerland four office rooms in a university research lab radiant heating with commercial electric heater with hot water radiators disabled $<1$ frequency regulation evaluation improved frequency regulation vs. simulated baseline
@@
MPC Berkeley, CA, US thermally identical test cells in a research facility forced-air with variable speed and dedicated AHUs 1 frequency regulation evaluation improved frequency regulation vs. measured baseline
@@
MPC Lausanne, Switzerland five offices in an office building electric heaters with direct modulation 1 frequency regulation evaluation improved frequency regulation vs. simulated baseline
@@
Granada, Spain three-story school hydronic heating with biomass boiler serving multiple circuits 43 minimize energy and improve comfort 33% energy savings and comfort improvement vs. measured baseline
@@
MPC Neuchâtel and Winterthur, Switzerland conference room and office building forced-air with AHUs, radiators, and rotating heat exchanger 10–38 minimize cost 21–23% cost savings vs. measured baseline
@@
MPC Halifax, NS, Canada five-story university building forced-air with VAV reheat, HPs, district heating, and rooftop cooling tower 113 minimize energy 29–63% energy savings vs. measured baseline
@@
MPC Salzburg, Austria two floors of a five-story university office building hybrid system with TABS (radiant) and fan coils (forced-air) 49 minimize energy experimental savings unreported; simulated energy savings 31–36% vs. simulated baseline
@@
MPC Sichuan Province, China six above-ground floors of a shopping mall forced-air with chillers, variable speed drivers, AHUs, and fan coils 1 minimize energy 16% energy savings vs. measured baseline
@@
RL Aachen, Germany three conference rooms in a university research building hybrid forced-air and radiant system with concrete core and air chillers 5 minimize energy 34.7% energy savings vs. measured baseline